# Probing the Bounce Energy Scale in Bouncing Cosmologies with Pulsar Timing Arrays

### v1

@@ III.1 Posterior distribution of bounce energy scale and physics interpretation
Figure 6: Violin plots of posterior SGWB amplitudes in each PTA frequency bin for two representative parameter sets—one from each of the $1\sigma$ posterior branches in the $(\rho_{s\downarrow}^{1/4},w_{1})$ plane (solid green and blue; Fig. 4), with examples in the $(a_{s\downarrow},w_{1})$ parametrization shown in solid purple and orange (Fig. 8). Dashed line correspond to six conventional SGWB models (Sec. A). The gray band shows the SMBHB astrophysical prediction [108]. Violins are built from the combined PTA datasets [2, 4, 6, 7, 99, 100, 101, 102, 103, 104, 105, 106] using the method of Ref. [32].
Furthermore, from a theoretical perspective the “left” branch at $w_{1}\approx 0.3$ satisfies all standard energy conditions—null ($w_{1}\geq-1$), weak ($\rho_{s\downarrow}\geq 0$ and $w_{1}\geq-1$), strong ($w_{1}\geq-\tfrac{1}{3}$), and dominant ($\rho_{s\downarrow}\geq 0$ and $-1\leq w_{1}\leq 1$)—as confirmed by the high-resolution inset in Fig. 5. By contrast, the “right” branch at $w_{1}\gg 1$ violates the dominant energy condition (Table 1). Violations of energy conditions have attracted considerable interest as effective descriptions of novel physics—e.g. ghost condensates, higher-derivative or modified-gravity operators, and extra-dimensional effects—operating during the very early Universe (inflation or the contracting phase of a bounce) (see, e.g., Refs. [45, 109, 110, 111, 112, 113, 114, 115, 98, 116, 117, 118]). Our PTA-based constraints therefore offer complementary empirical motivation for pursuing bounce models that violate the dominant energy condition.
Table 1: Best‐fit values for $n_{T}$, $w_{1}$ and $\rho_{s\downarrow}^{1/4}$ determined in Bayesian inference
Dataset $w_{1}$ range Best‐fit $n_{T}$ Best‐fit $w_{1}$ Best‐fit $\log_{10}(\rho_{s}^{1/4}/\mathrm{TeV})$
NG 15 $[-\tfrac{1}{3},1]$ $1.871$ $0.293$ 31.317
EPTA DR2 $[-\tfrac{1}{3},1]$ $2.120$ $0.376$ 35.885
PPTA DR3 $[-\tfrac{1}{3},1]$ $1.676$ $0.240$ 28.053
IPTA DR2 $[-\tfrac{1}{3},1]$ $2.262$ $0.434$ 39.147
@@
Another point of theoretical interest is that the best‐fit bounce energy scale exceeds the Planck scale (Table 1),
@@
for both posterior branches. This finding is consistent with the PTA‐only lower bound derived in Ref. [97], where SGWB limits and forecasts—from PTAs, Planck/BICEP, aLIGO+aVirgo (O2), CMB-S4, IPTA (design), SKA, DECIGO, BBO, TianQin, Taiji, aLIGO+aVirgo+KAGRA (design), Cosmic Explorer, and Einstein Telescope—were used to constrain $\rho_{s\downarrow}^{1/4}$ across multiple bands [119, 120, 121].Hence, current pulsar-timing arrays are already probing trans-Planckian physics in bouncing-cosmology scenarios. Any viable UV completion must therefore respect these PTA bounds, providing a rare observational window into physics beyond the Planck scale. This situation echoes the “trans-Planckian problem” of inflation [122, 123, 124, 125, 126, 127, 128, 129], in which modes originating above $m_{\rm pl}$ nonetheless leave imprints on low-energy observables, highlighting the need for new quantum-gravitational input to render predictions unambiguous.
III.2 Bayes‐factor comparison with conventional SGWB sources
@@
Table 2: Bayes factors for the bouncing‐universe model $(\rho_{s\downarrow}^{1/4},w_{1})$ versus six conventional SGWB source models.
@@
Table 3: Interpretation of Bayes‐factor values for model comparison, following the Jeffreys scale [12, 130].
@@
III.3 Comparing with the dual scenario of inflation-bounce cosmologies
The dual scenario of inflation-bounce cosmologies $(w,r)$ [94], obtained by extending Wands’ duality [131, 132, 133, 134], predicts both a nearly scale‐invariant CMB spectrum $(n_{s}-1\simeq 0)$ and a blue‐tilted SGWB $(n_{T}=1.8\pm 0.3)$ consistent with CMB/BICEP and PTA data [45] (Fig. 7). Our bouncing‐universe model is a concrete realization of this scenario, with an explicit SGWB spectrum $\Omega_{\mathrm{GW}}(f)=\Omega_{\mathrm{GW}}\bigl{(}f,\rho_{s\downarrow}^{1/4}\bigr{)}$ that introduces the bounce energy scale $\rho_{s\downarrow}^{1/4}$. Because the form of $\Omega_{\mathrm{GW}}(f,\rho_{s\downarrow}^{1/4})$ is fixed in this concrete realization, PTAs impose tighter constraints on the $(\rho_{s\downarrow}^{1/4},w_{1})$ parameter space, yielding smaller Bayes factors relative to the more generic $(w,r)$ dual scenario. Table 4 presents the Bayes‐factor comparison between $(\rho_{s\downarrow}^{1/4},w_{1})$ and $(w,r)$, confirming that PTA data favor the concrete bouncing‐universe realization while still supporting the overarching dual scenario.More specifically, Table 4 shows that every Bayes factor for $(\rho_{s\downarrow}^{1/4},w_{1})$ versus either branch—or the bounce‐only branch—of the dual scenario $(w,r)$ lies below unity, as anticipated.
Figure 7: Posterior distribution of dual scenario of inflation-bounce cosmologies $(w,r)$, obtained by fitting its SGWB spectrum to the same full PTA dataset, c.f. Figure 3 in Ref. [45].
Table 4: Bayes factors for the bouncing‐universe model $(\rho_{s\downarrow}^{1/4},w_{1})$ versus dual scenario of inflation-bounce cosmologies $(w,r)$ [45, 94].
@@ III.3 Comparing with the dual scenario of inflation-bounce cosmologies
IV Posterior distribution of bounce scale factor
In this section we recast the PTA bounds on the bounce energy scale $\rho_{s\downarrow}^{1/4}$ into an equivalent constraint on the bounce scale factor $a_{s\downarrow}$. Here $a_{s\downarrow}$ is defined as the quasi‐minimal size of the Universe at the bounce (the boundary between Phases III and IV), satisfying
@@ IV Posterior distribution of bounce scale factor
$\Omega_{\rm GW}(f)\,h^{2}=\frac{h^{2}}{24}\,\Bigl{(}\frac{f_{H_{0}}}{f_{m_{\rm pl}}}\Bigr{)}^{2}\,\frac{C(w_{1})}{(2\pi)^{-n_{T}-1}}\,\Bigl{(}\frac{f}{f_{H_{0}}}\Bigr{)}^{n_{T}}\Bigl{(}\frac{a_{s\downarrow}}{\sqrt{\Omega_{\gamma 0}}}\Bigr{)}^{-(4-n_{T})}\,\mathcal{T}_{\rm eq}(f)\,,$ (10)
@@
By performing a Bayesian fit of Eqs. (10) and (1) to the full PTA datasets (NANOGrav 15-yr, EPTA DR2, PPTA DR3, IPTA DR2) [2, 4, 6, 7, 99, 100, 101, 102, 103, 104, 105, 106], we derive the two-dimensional posterior in the $(a_{s\downarrow},w_{1})$ plane for a generic bouncing cosmology (Fig. 8). These constraints on the bounce scale factor $a_{s\downarrow}$ are equivalent—via Eq. (9)—to those on $\rho_{s\downarrow}^{1/4}$ shown in Fig. 4, providing a dual characterization of the bounce energy scale.Akin to the constraints on $\rho_{s\downarrow}^{1/4}$, the $(a_{s\downarrow},w_{1})$ posterior (Fig. 8) exhibits two notable features with implications for novel physics: (1) The “left” branch satisfies all standard energy conditions, whereas the “right” branch violates the dominant energy condition, thereby offering complementary empirical motivation to explore mechanisms—such as ghost condensates or modified-gravity effects—that permit energy‐condition violation during the very early Universe. (2) The best‐fit bounce scale factor $a_{s\downarrow}$ lies below the Planck scale factor
@@
In Table 5 we present the Bayes factors for the bouncing‐universe model $(a_{s\downarrow},w_{1})$ versus six conventional SGWB sources. Nearly all values exceed unity, indicating a generic preference for the bounce scenario. These findings agree with the Bayes‐factor results for $(\rho_{s\downarrow}^{1/4},w_{1})$ in Table 2; the minor differences arise from the chosen prior range in $a_{s\downarrow}$ used in the inference.
Table 5: Bayes factors for the bouncing‐universe model $(a_{s\downarrow},w_{1})$ versus six conventional SGWB sources
@@
V Summary and Prospects
We have presented the first Bayesian determination of the bounce energy scale $\rho_{s\downarrow}^{1/4}$ in a generic bouncing-universe model that can account for the nanohertz SGWB measured by PTAs (NANOGrav 15-yr, EPTA DR2, PPTA DR3, IPTA DR2). Fitting the analytic SGWB spectrum derived in Ref. [96], we uncover two viable posterior branches in $(\rho_{s\downarrow}^{1/4},w_{1})$: one near $w_{1}\approx 0.3$ and one at $w_{1}\gg 1$.
@@ V Summary and Prospects
Finally, we have shown (Sec. B) that a normalizing-flow–based machine-learning emulator reproduces our MCMC posterior at greatly reduced cost, paving the way for rapid Bayesian inference as PTA datasets—especially with future facilities such as SKA—continue to expand.
@@ A.1 Model descriptions
In this section we summarize each SGWB source model, following the definitions and notation of Refs. [45, 130].
@@
    Model-3: SMBHBs (supermassive-black-hole binaries) [135, 27, 136, 137]
@@
    Model-4: Inflationary GWs/dual scenario $(n_{T},r)$ [14, 45]
@@
    Model-5: Cosmic strings (CS) [9, 138, 139, 140, 141]
@@
    Model-6: Domain walls (DW) [25, 142, 143]
    $\Omega_{\rm GW}^{\rm DW}h^{2}=\Omega_{\mathrm{GW}}^{\text{peak}}h^{2}\,S^{\rm dw}(f),\quad S^{\rm dw}(f)=\begin{cases}(f/f_{\rm peak})^{3},&f<f_{\rm peak},\\ (f/f_{\rm peak})^{-1},&f\geq f_{\rm peak},\end{cases}$ (15)
    with
    $\Omega_{\mathrm{GW}}^{\text{peak}}h^{2}\simeq 5.20\times 10^{-20}\times\tilde{\epsilon}_{\mathrm{gw}}\mathcal{A}^{4}\left(\frac{10.75}{g_{\ast}}\right)^{1/3}\times\left(\frac{\sigma}{1\mathrm{TeV}^{3}}\right)^{4}\left(\frac{1\mathrm{MeV}^{4}}{\Delta V}\right)^{2}$ (16)
    where $\Omega_{\mathrm{GW}}^{\text{peak}}$ is peak SGWB spectrum; $f_{\rm peak}$ peak frequency; $\sigma$ wall tension; $\Delta V$ vacuum bias; $\mathcal{A}=1.2$ DW area parameter, and $\tilde{\epsilon}_{\rm gw}=0.7$ efficiency factor.
@@
    Model-7: First-order phase transition (FOPT) [30, 12, 144, 143]
    $\Omega_{\rm GW}^{\rm FOPT}h^{2}=2.65\times 10^{-6}(H_{\ast}\tau_{\rm sw})(\beta/H_{\ast})^{-1}v_{b}\Bigl{(}\frac{\kappa_{v}\alpha_{PT}}{1+\alpha_{PT}}\Bigr{)}^{2}F(f/f_{\text{peak}}^{\mathrm{FOPT}}),$ (17)
@@
    Model-8: Scalar-induced GWs (SIGW) [36, 12]
[…]
@@
    $\mathcal{K}(u,v)$ $=\left(\frac{4v^{2}-\left(1+v^{2}-u^{2}\right)^{2}}{4uv}\right)^{2}\times\frac{1}{2}\left(\frac{3}{4u^{3}v^{3}x}\right)^{2}\left(u^{2}+v^{2}-3\right)^{2}$ (20)
@@
    $\left.+\left[\pi\left(u^{2}+v^{2}-3\right)\Theta(u+v-\sqrt{3})\right]^{2}\right\}.$
@@
    Model-9: Dual scenario (inflation+bounce) $(w,r)$ [45]
@@
    Model-10: Dual scenario (only bounce) $(w>1,r)$ [45]. Same as Model-9 with prior restricted to $w>1/3$.

### v4

@@ 3.1 Posterior distribution of bounce energy scale and physics interpretation
Figure 6: Violin plots of posterior SGWB amplitudes in each PTA frequency bin for two representative parameter sets—one from each of the $1\sigma$ posterior branches in the $(\rho_{s\downarrow}^{1/4},w_{1})$ plane (solid green and blue; Fig. 4), with examples in the $(a_{s\downarrow},w_{1})$ parametrization shown in solid purple and orange (Fig. 8). Dashed line correspond to six conventional SGWB models (Sec. A). The gray band shows the SMBHB astrophysical prediction Rosado:2015epa . Violins are built from the combined PTA datasets NANOGrav:2023gor ; EPTA:2023fyk ; Reardon:2023gzh ; Antoniadis:2022pcn ; ZenodoNG ; NANOGrav:2023hde ; ZenodoEPTA ; EPTA:2023sfo ; PPTADR3 ; Zic:2023gta ; IPTADR2 ; Perera:2019sca using the method of Ref. Gouttenoire:2023bqy .
Furthermore, from a theoretical perspective the “left” branch at $w_{1}\approx 0.3$ satisfies all standard energy conditions—null ($w_{1}\geq-1$), weak ($\rho_{s\downarrow}\geq 0$ and $w_{1}\geq-1$), strong ($w_{1}\geq-\tfrac{1}{3}$), and dominant ($\rho_{s\downarrow}\geq 0$ and $-1\leq w_{1}\leq 1$)—as confirmed by the high-resolution inset in Fig. 5. By contrast, the “right” branch at $w_{1}\gg 1$ violates the dominant energy condition (Table 1). Violations of energy conditions have attracted considerable interest as effective descriptions of novel physics—e.g. ghost condensates, higher-derivative or modified-gravity operators, and extra-dimensional effects—operating during the very early Universe (inflation or the contracting phase of a bounce) (see, e.g., Refs. Li:2024oru ; Cai:2023uhc ; Ye:2023tpz ; Lohakare:2022umj ; Gouttenoire:2021jhk ; Bhattacharya:2015nda ; Cai:2013kja ; Liu:2012fk ; Cai:2009zp ; Santos:2007pp ; Santos:2007bs ; Burgess:2003tz ). Our PTA-based constraints therefore offer complementary empirical motivation for pursuing bounce models that violate the dominant energy condition.
Table 1: Best-fit values for $n_{T}$, $w_{1}$ and $\rho_{s\downarrow}^{1/4}$ determined in Bayesian inference
Dataset $w_{1}$ range Best-fit $n_{T}$ Best-fit $w_{1}$ Best-fit $\log_{10}(\rho_{s}^{1/4}/\mathrm{TeV})$
NG 15 $[-\frac{1}{3},1]$ $1.871$ $0.293$ 31.317
EPTA DR2 $[-\frac{1}{3},1]$ $2.120$ $0.376$ 35.885
PPTA DR3 $[-\frac{1}{3},1]$ $1.676$ $0.240$ 28.053
IPTA DR2 $[-\frac{1}{3},1]$ $2.262$ $0.434$ 39.147
@@
Another point of theoretical interest is that the best-fit bounce energy scale exceeds the Planck scale (Table 1),
@@
for both posterior branches. This finding is consistent with the PTA-only lower bound derived in Ref. Li:2025ilc , where SGWB limits and forecasts—from PTAs, Planck/BICEP, aLIGO+aVirgo (O2), CMB-S4, IPTA (design), SKA, DECIGO, BBO, TianQin, Taiji, aLIGO+aVirgo+KAGRA (design), Cosmic Explorer, and Einstein Telescope—were used to constrain $\rho_{s\downarrow}^{1/4}$ across multiple bands Schmitz:2020syl ; Annis:2022xgg ; Bi:2023tib .Hence, current pulsar-timing arrays are already probing trans-Planckian physics in bouncing-cosmology scenarios. Any viable UV completion must therefore respect these PTA bounds, providing a rare observational window into physics beyond the Planck scale. This situation echoes the “trans-Planckian problem” of inflation Martin:2000xs ; Brandenberger:2000wr ; Brandenberger:2012aj ; Kaloper:2002cs ; Easther:2002xe ; Bedroya:2019snp ; Bedroya:2019tba ; Benetti:2021uea ; Bedroya:2022tbh (see  Bedroya:2024zta ; Bedroya:2025ris for recent development.), in which modes originating above $m_{\rm pl}$ nonetheless leave imprints on low-energy observables, highlighting the need for new quantum-gravitational input to render predictions unambiguous.
3.2 Bayes-factor comparison with conventional SGWB sources
@@
Table 2: Bayes factors for the bouncing-universe model $(\rho_{s\downarrow}^{1/4},w_{1})$ versus six conventional SGWB source models.
@@
Table 3: Interpretation of Bayes-factor values for model comparison, following the Jeffreys scale Bian:2023dnv ; Lai:2025xov .
@@
3.3 Comparing with the dual scenario of inflation-bounce cosmologies
The dual scenario of inflation-bounce cosmologies $(w,r)$ Li:2013bha , obtained by extending Wands’ duality Wands:1998yp ; Finelli:2001sr ; Boyle:2004gv ; Raveendran:2023auh , predicts both a nearly scale-invariant CMB spectrum $(n_{s}-1\simeq 0)$ and a blue-tilted SGWB $(n_{T}=1.8\pm 0.3)$ consistent with CMB/BICEP and PTA data Li:2024oru (Fig. 7). Our bouncing-universe model is a concrete realization of this scenario, with an explicit SGWB spectrum $\Omega_{\mathrm{GW}}(f)=\Omega_{\mathrm{GW}}\bigl{(}f,\rho_{s\downarrow}^{1/4}\bigr{)}$ that introduces the bounce energy scale $\rho_{s\downarrow}^{1/4}$. Because the form of $\Omega_{\mathrm{GW}}(f,\rho_{s\downarrow}^{1/4})$ is fixed in this concrete realization, PTAs impose tighter constraints on the $(\rho_{s\downarrow}^{1/4},w_{1})$ parameter space, yielding smaller Bayes factors relative to the more generic $(w,r)$ dual scenario. Table 4 presents the Bayes-factor comparison between $(\rho_{s\downarrow}^{1/4},w_{1})$ and $(w,r)$, confirming that PTA data favor the concrete bouncing-universe realization while still supporting the overarching dual scenario.More specifically, Table 4 shows that every Bayes factor for $(\rho_{s\downarrow}^{1/4},w_{1})$ versus either branch—or the bounce-only branch—of the dual scenario $(w,r)$ lies below unity, as anticipated.
Figure 7: Posterior distribution of dual scenario of inflation-bounce cosmologies $(w,r)$, obtained by fitting its SGWB spectrum to the same full PTA dataset, c.f. Figure 3 in Ref. Li:2024oru .
Table 4: Bayes factors for the bouncing-universe model $(\rho_{s\downarrow}^{1/4},w_{1})$ versus dual scenario of inflation-bounce cosmologies $(w,r)$ Li:2024oru ; Li:2013bha .
@@ 3.3 Comparing with the dual scenario of inflation-bounce cosmologies
4 Posterior distribution of bounce scale factor
In this section we recast the PTA bounds on the bounce energy scale $\rho_{s\downarrow}^{1/4}$ into an equivalent constraint on the bounce scale factor $a_{s\downarrow}$. Here $a_{s\downarrow}$ is defined as the quasi-minimal size of the Universe at the bounce (the boundary between Phases III and IV), satisfying
@@ 4 Posterior distribution of bounce scale factor
$\begin{split}\Omega_{\rm GW}(f)\,h^{2}&=\frac{h^{2}}{24}\,\Bigl{(}\frac{f_{H_{0}}}{f_{m_{\rm pl}}}\Bigr{)}^{2}\,\frac{C(w_{1})}{(2\pi)^{-n_{T}-1}}\,\Bigl{(}\frac{f}{f_{H_{0}}}\Bigr{)}^{n_{T}}\\ &\quad\times\Bigl{(}\frac{a_{s\downarrow}}{\sqrt{\Omega_{\gamma 0}}}\Bigr{)}^{-(4-n_{T})}\,\mathcal{T}_{\rm eq}(f)\,,\end{split}$ (10)
@@
By performing a Bayesian fit of Eqs. (10) and (1) to the full PTA datasets (NANOGrav 15-yr, EPTA DR2, PPTA DR3, IPTA DR2) NANOGrav:2023gor ; EPTA:2023fyk ; Reardon:2023gzh ; Antoniadis:2022pcn ; ZenodoNG ; NANOGrav:2023hde ; ZenodoEPTA ; EPTA:2023sfo ; PPTADR3 ; Zic:2023gta ; IPTADR2 ; Perera:2019sca , we derive the two-dimensional posterior in the $(a_{s\downarrow},w_{1})$ plane for a generic bouncing cosmology (Fig. 8). These constraints on the bounce scale factor $a_{s\downarrow}$ are equivalent—via Eq. (9)—to those on $\rho_{s\downarrow}^{1/4}$ shown in Fig. 4, providing a dual characterization of the bounce energy scale.Akin to the constraints on $\rho_{s\downarrow}^{1/4}$, the $(a_{s\downarrow},w_{1})$ posterior (Fig. 8) exhibits two notable features with implications for novel physics: (1) The “left” branch satisfies all standard energy conditions, whereas the “right” branch violates the dominant energy condition, thereby offering complementary empirical motivation to explore mechanisms—such as ghost condensates or modified-gravity effects—that permit energy-condition violation during the very early Universe. (2) The best-fit bounce scale factor $a_{s\downarrow}$ lies below the Planck scale factor
@@
In Table 5 we present the Bayes factors for the bouncing-universe model $(a_{s\downarrow},w_{1})$ versus six conventional SGWB sources. Nearly all values exceed unity, indicating a generic preference for the bounce scenario. These findings agree with the Bayes-factor results for $(\rho_{s\downarrow}^{1/4},w_{1})$ in Table 2; the minor differences arise from the chosen prior range in $a_{s\downarrow}$ used in the inference.
Table 5: Bayes factors for the bouncing-universe model $(a_{s\downarrow},w_{1})$ versus six conventional SGWB sources
@@
5 Summary and Prospects
We have presented the first Bayesian determination of the bounce energy scale $\rho_{s\downarrow}^{1/4}$ in a generic bouncing-universe model that can account for the common-spectrum process, possibly the first hint of a nanohertz SGWB, reported by PTAs (NANOGrav 15-yr, EPTA DR2, PPTA DR3, IPTA DR2). Fitting the analytic SGWB spectrum derived in Ref. Li:2024dce , we uncover two viable posterior branches in $(\rho_{s\downarrow}^{1/4},w_{1})$: one near $w_{1}\approx 0.3$ and one at $w_{1}\gg 1$.
@@ 5 Summary and Prospects
Finally, we have shown (Sec. B) that a normalizing-flow–based machine-learning emulator reproduces our MCMC posterior at greatly reduced cost, paving the way for rapid Bayesian inference as PTA datasets—especially with future facilities such as SKA—continue to expand. Last but not least, in Sec. C we discuss the detectability of the bouncing-universe SGWB by space- and ground-based GW searches at higher frequencies, as well as additional constraints from BBN and CMB bounds. [^1] We thank the anonymous referee for this suggestion, which has illuminated directions for future work. These considerations imply that a broken power-law feature must be incorporated into the SGWB spectrum of bouncing cosmologies to satisfy BBN and CMB limits and to enable detection by current and forthcoming space- and ground-based GW detectors at higher frequencies.
@@ A.1 Model descriptions
In this section we summarize each SGWB source model, following the definitions and notation of Refs. Li:2024oru ; Lai:2025xov .
@@
    Model-3: SMBHBs (supermassive-black-hole binaries) Hobbs:2009yn ; Bian:2022qbh ; Jaffe:2002rt ; Wyithe:2002ep
@@
    Model-4: Inflationary GWs/dual scenario $(n_{T},r)$ Caprini:2018mtu ; Li:2024oru
@@
    Model-5: Cosmic strings (CS) NANOGrav:2023hvm ; Blanco-Pillado:2017oxo ; Chang:2021afa ; Auclair:2019wcv ; Martins:2000cs
@@
    Model-6: Domain walls (DW) Hiramatsu:2013qaa ; Kadota:2015dza ; Zhou:2020ojf
    $\Omega_{\rm GW}^{\rm DW}h^{2}$ $=\Omega_{\mathrm{GW}}^{\text{peak}}h^{2}\,S^{\rm dw}(f),$ (15a)
    $S^{\rm dw}(f)$ =$=\begin{cases}\bigl{(}f/f_{\rm peak}\bigr{)}^{3},&f<f_{\rm peak},\\ \bigl{(}f/f_{\rm peak}\bigr{)}^{-1},&f\geq f_{\rm peak},\end{cases}$ (15b)
    $\begin{split}\Omega_{\mathrm{GW}}^{\text{peak}}h^{2}&\simeq 5.20\times 10^{-20}\,\tilde{\epsilon}_{\mathrm{gw}}\,\mathcal{A}^{4}\Bigl{(}\frac{10.75}{g_{\ast}}\Bigr{)}^{1/3}\\ &\quad\times\Bigl{(}\frac{\sigma}{1\,\mathrm{TeV}^{3}}\Bigr{)}^{4}\Bigl{(}\frac{1\,\mathrm{MeV}^{4}}{\Delta V}\Bigr{)}^{2}\end{split}$ (16)
    where $\Omega_{\mathrm{GW}}^{\text{peak}}$ is peak SGWB spectrum; $f_{\rm peak}$ peak frequency; $\sigma$ wall tension; $\Delta V$ vacuum bias; $\mathcal{A}=1.2$ DW area parameter, and $\tilde{\epsilon}_{\rm gw}=0.7$ efficiency factor. For recent development, see Refs. Babichev:2023pbf ; Dankovsky:2024ipq
@@
    Model-7: First-order phase transition (FOPT) Caprini:2015zlo ; Bian:2023dnv ; Hirose:2023bvl ; Zhou:2020ojf
    $\begin{split}\Omega_{\rm GW}^{\rm FOPT}h^{2}&=2.65\times 10^{-6}\,(H_{\ast}\tau_{\rm sw})\,(\beta/H_{\ast})^{-1}v_{b}\\ &\quad\times\Bigl{(}\frac{\kappa_{v}\alpha_{PT}}{1+\alpha_{PT}}\Bigr{)}^{2}F\bigl{(}f/f_{\text{peak}}^{\mathrm{FOPT}}\bigr{)}\,,\end{split}$ (17)
@@
    Model-8: Scalar-induced GWs (SIGW) Kohri:2018awv ; Bian:2023dnv
    $\Omega_{\rm GW}^{\rm SI}h^{2}$ $=\tfrac{1}{12}\,\Omega_{\rm rad}h^{2}\Bigl{(}\tfrac{g_{0}}{g_{\ast}}\Bigr{)}^{1/3}$ (19)
    ×$\qquad\times\int\!du\,dv\,\mathcal{K}(u,v)\,P_{\mathcal{R}}(ku)\,P_{\mathcal{R}}(kv)\,.$
@@
    $\mathcal{K}(u,v)$ $=\left(\frac{4v^{2}-\bigl{(}1+v^{2}-u^{2}\bigr{)}^{2}}{4uv}\right)^{2}$ (20)
    ×$\times\frac{1}{2}\!\left(\frac{3}{4u^{3}v^{3}x}\right)^{2}\,(u^{2}+v^{2}-3)^{2}$
@@
    $\left.+\left[\pi\left(u^{2}+v^{2}-3\right)\Theta(u+v-\sqrt{3})\right]^{2}\right\},$
@@
    Model-9: Dual scenario (inflation+bounce) $(w,r)$ Li:2024oru
@@
    Model-10: Dual scenario (only bounce) $(w>1,r)$ Li:2024oru . Same as Model-9 with prior restricted to $w>1/3$.